I have to add 15ml of fertilizer to 50 litres of water, how much fertilizer is that per litre. Thank you.
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

15/50=3/10 ml per litre

Answer by Alan3354(69443)

I have to add 15ml of fertilizer to 50 litres of water, how much fertilizer is that per litre.
--------------
The total is 15 + 50000 = 50015 ml
15/50015 = 3/10003 ml of fertilizer per liter
=~ 0.29991 ml/liter, or practically 0.3 ml/liter